# TMT: Tri-Modal Translation between Speech, Image, and Text by Processing Different Modalities as Different Languages

### v1

@@ 4.2 Metrics
Captioning tasks. For the image captioning and image-to-speech captioning tasks, we employ BLEU Papineni et al. (2002), METEOR Denkowski and Lavie (2014), ROUGE Lin (2004), CIDEr Vedantam et al. (2015), and SPICE Anderson et al. (2016). BLEU, METEOR, ROUGE, and CIDEr measure n-gram overlaps between predictions and references. SPICE compares semantic propositional content. Across all metrics, higher value indicates better performance.
Image synthesis tasks. For the text- and speech-driven image synthesis tasks, we adopt the CLIP score Radford et al. (2021) (i.e., the cosine similarity) to assess how similar the generated image is associated with the input text or speech, following Ge et al. (2023a); Koh et al. (2023).
ASR and TTS. We employ the widely used Word Error Rate (WER) for ASR and TTS. For TTS, we transcribe the generated speech by using a pre-trained ASR model Baevski et al. (2020), following Lakhotia et al. (2021); Kim et al. (2023). Additionally, we employ Mean Opinion Score (MOS) and neural MOS to assess the quality of generated results.
@@
Raw Image SAT Xu et al. (2015) 1 24.3 23.9 - - - 21.3 20.3 - - -
BLIP-2 Li et al. (2023a) 1 43.7 - - 145.8 - - - - - -
Discrete Image Token Kim et al. (2024a) 1 29.9 25.2 52.8 97.4 18.6 23.4 22.0 48.9 63.3 15.4
Ge et al. (2023a) 2 34.6 28.4 56.4 131.0 22.9 - - - - -
@@
Table 4: Image captioning performance (Image-to-Text Translation) comparisons on COCO and Flickr8k. We also report the performance of raw image-based models for reference purposes. “Num. MMT.” indicates the number of MMT tasks that the model is capable of.
@@
Raw Image Wang et al. (2021) 1 - - - - - 3.5 11.3 23.2 8.0 -
Hsu et al. (2021b) 1 23.3 21.2 47.8 73.2 14.9 12.5 14.5 39.1 24.5 9.5
Effendi et al. (2021) 1 - - - - - 14.8 17.4 32.9 45.8 -
Kim et al. (2024a) 1 25.9 23.8 50.4 81.1 17.5 20.6 22.0 48.4 53.6 15.8
Discrete Image Token Kim et al. (2024a) 1 20.1 21.4 46.4 64.0 15.0 16.7 19.6 44.2 41.2 13.1
@@
Table 5: Image-to-speech captioning performance (Image-to-Speech Translation) comparisons on COCO and Flickr8k. We also report the performance of raw image-based models for reference purposes.
4.3 Implementation Details
The input image is resized into 224 $\times$ 224, while the output image size is 768 $\times$ 768. The input audio is resampled to 16kHz and tokenized using a pre-trained HuBERT base model Hsu et al. (2021a). The backbone model that models the multi-modal information is an encoder-decoder Transformer Vaswani et al. (2017) with 12 layers each. For pre-training the TMT before applying BT, we train the model for 350k steps. Then, we further train the model using the additional data obtained by applying BT, for 250k steps with the same configuration but with a peak learning rate of $5e^{-5}$. For decoding, we use beam search with a beam width of 5. Full implementation details including hyper-parameters can be found in appendix B.
4.4 Results
Firstly, we illustrate selected samples for each task of TMT in Fig. 2. We then compare single MMT task models versus the proposed TMT model. Next, we compare our TMT with other works in the literature for each MMT task. Finally, we conduct a human subjective evaluation on two speech generation tasks: Image-to-Speech synthesis and TTS. Further ablations on the discrete token vocabulary size and image tokenizer are presented in appendix C.
4.4.1 Effectiveness of Unifying Tasks
@@
Table 2 compares the captioning performances, including image-to-text and image-to-speech translations. The results show that by unifying tasks, we can consistently obtain performance gain. In our analysis, this stems from the unified multi-modal decoder. The decoder can leverage complementary multi-modal information in learning language modeling. Table 3 shows the comparison results for the remaining four tasks. Similar to captioning, the unified TMT model outperforms its single-task counterparts consistently, except for the text-to-image translation. However, the performance differences between single MMT and TMT in text-to-image translation are marginal, while the performance gain obtained by unifying tasks is substantial for speech-to-image translation. We conclude that unifying tasks is beneficial not only for practicality by building a single model but also in the aspect of performance.
4.4.2 Comparison with the literature
We compare TMT’s performance on each MMT task with the other works in the literature. The TMT model in this section employs 10M training samples.
@@ 4.4.2 Comparison with the literature
Image Captioning Tasks. Table 4 shows the results of image-to-text translation of TMT and other competing systems. We also report the performances of raw image-based systems for reference purposes. We observe that our TMT accomplishes comparable performances among state-of-the-art discrete image-based models. Notably, the proposed TMT is competitive with the LLM-based method Ge et al. (2023a) with a much smaller number of parameters, on COCO. Note that Ge et al. (2023a) is built on OPT 2.7B model Zhang et al. (2022) while the proposed TMT consists of a 270M multi-modal model. This shows the effectiveness of the proposed TMT compared to the method of finetuning LLM. Moreover, TMT can perform six MMT tasks while the previous methods can only perform two MMT tasks at maximum.
@@
Table 5 shows the direct image-to-speech captioning performances. TMT outperforms the recently proposed state-of-the-art Kim et al. (2024a) on both COCO and Flickr8k databases and even achieves comparable performances with raw-image based method. Since the decoder of the proposed TMT is trained using multi-tasks with text simultaneously, it better models linguistics compared to the transfer learning from the text of Kim et al. (2024a).
@@
Image Synthesis Tasks. The upper part of Table 6 displays the performances of text- and speech-driven image synthesis tasks. Following the literature, we evaluate the model on Flickr30k instead of using Flickr8k for the text-to-image synthesis tasks. The proposed TMT achieves the best CLIP similarity on both databases in text-to-image translation, indicating that TMT generates images that best describe the given text sentences. Note that the proposed TMT has much fewer model parameters (i.e., 270M) compared to Koh et al. (2023) and Ge et al. (2023a) which are based on LLMs, OPT 6.7B, and OPT 2.7B, respectively. TMT still outperforms the competitors with more diverse MMT task abilities. Also, the speech-to-image translation performance of TMT on COCO is 68.88, closer to that of text-to-image performance. We hence confirm that even if using speech directly as input, TMT correctly generates images described in the speech. We underline that this work is the first to explore high-resolution image synthesis at a resolution of 768$\times$768 from speech. In addition to the samples shown in Fig 2, appendix A.
@@
ASR and TTS Tasks. The lower part of Table 6 shows the performances of ASR and TTS. For the evaluation purpose, we compare the ASR performance with a pre-trained Wav2Vec2.0 (W2V2) Baevski et al. (2020) model. The results show that TMT can correctly recognize the input speech into text by achieving lower WER compared to W2V2. For TTS, as the proposed model is a discrete token-based speech synthesis model, we compare the performance with the resynthesized speech (ReSyn) from the ground-truth speech tokens following Lakhotia et al. (2021); Kim et al. (2023). Results demonstrate that the TMT can generate the correct speech from input text without suffering in synthesizing background noises which inherently exist in the ground-truth speech tokens.
@@
Table 6: Text- and Speech-driven image synthesis, ASR, and TTS (Text-to-Image, Speech-to-Image, Speech-to-Text, and Text-to-Speech Translation) performance comparisons on COCO and Flickr8k. For TTS, we report the WER by feeding the synthesized audio into a pre-trained ASR model following Kim et al. (2023).
4.4.3 Mean Opinion Score Comparisons
In order to assess the quality of generated speech in both image-to-speech captioning and text-to-speech tasks, we conduct human subjective study and neural MOS test using MOSNet Lo et al. (2018) and SpeechLMScore Maiti et al. (2023). Specifically, we enlisted 21 participants to evaluate the naturalness of the generated speech for both tasks, and for the image-to-speech captioning task, we also assessed descriptiveness, gauging the extent to which the speech accurately describes the input images. Following the previous works Hsu et al. (2021b); Kim et al. (2024a), 20 samples are used for the image-to-speech captioning task. For TTS, 20 samples and 100 samples are employed for human subjective study and neural MOS, respectively. The results are shown in Table 7. TMT achieves the best score for both human study and neural MOS scores in image-to-speech captioning. Moreover, we achieve comparable performances with VITS on TTS task. Through MOS evaluation, we can confirm the effectiveness of TMT in generating intelligible and natural speech.
  Methods MOS Neural MOS
Natural Descriptive MOSNet $\uparrow$ SpeechLM $\downarrow$
$\bullet$ Image-to-Speech Captioning
Hsu et al. (2021b) 2.70±${}_{\pm 0.38}$ 2.96±${}_{\pm 0.22}$ 2.51 4.23
Kim et al. (2024a) 3.71±${}_{\pm 0.25}$ 3.41±${}_{\pm 0.23}$ 4.30 4.16
[…]
$\bullet$ Text-to-Speech Synthesis (TTS)
VITS Kim et al. (2021a) 3.85±${}_{\pm 0.25}$ 4.64 4.02
[…]
Table 7: Human evaluation score (MOS) comparisons with 95% confidence interval, and Neural MOS scores in image-to-speech captioning and TTS on COCO.
  Methods Image Captioningnormal-$\uparrow$ Image-to-Speechnormal-$\uparrow$ ASRnormal-$\downarrow$
TMT w/o BT 56.7 48.9 5.24
TMT 56.9 49.5 5.15
  Text-to-Imagenormal-$\uparrow$ Speech-to-Imagenormal-$\uparrow$ TTSnormal-$\downarrow$
TMT w/o BT 69.48 68.67 9.78
TMT 69.57 68.88 9.54
Table 8: Effectiveness of training with additional data derived with Back Translation (BT) on COCO. ROUGE is reported for Image Captioning and Image-to-Speech Captioning tasks.
4.4.4 Effectiveness of Back Translation
Table 8 shows the result of training TMT with and without BT. By employing additional 4M back translated data, we observe no significant yet consistent improvement across tasks of TMT. This shows the potential for further improvement, since the amount of additional data BT can produce is limitless. In appendix D, we show the examples of back translated data on both uni-modal data and bi-modal data.
5 Conclusion
We introduced TMT, a novel tri-modal translation model between speech, image, and text modalities. We interpreted different modalities as new languages and approached the MMT task as an NMT task after tokenizing all modalities. Our experiments demonstrated that the three modalities could be successfully translated using a multi-modal encoder-decoder architecture, incorporating six MMT tasks in a single model. TMT outperformed single MMT model counterparts. Remarkably, TMT achieves comparable performances with LLM-based methods consisting of more than 2.7B parameters, whereas TMT only has 270M parameters.
Appendix A Additional Qualitative Results
Fig. 3 illustrates additional generated images by using the text and speech as inputs (i.e.Text-driven Image Synthesis and Speech-driven Image Synthesis).
Appendix B Implementation details
[…]
Appendix C Ablation on Speech and Image Tokens
@@
To confirm the effectiveness of the different multi-modal tokens, we perform ablation studies by differing configurations for speech and image tokens. First, we differentiate the number of k-means clustering of the speech tokenizer to obtain different token sizes. Then, we train single MMT models of image-to-speech captioning and ASR using 200, 500, and 1000 token sizes. Table 9 shows the image-to-speech captioning and ASR performances. The results show that there is a trade-off in performances of ASR (WER) and image-to-speech captioning (BLEU, METEOR, ROUGE, CIDEr, SPICE). We confirm a consistent tendency where image-to-speech captioning performs better with a smaller number of speech tokens. In contrast, for ASR, employing a large number of tokens is beneficial. Since the performance gaps in ASR are smaller than those of image-to-speech captioning, we employ 200 speech tokens for all experiments.
Second, we investigate the different types of visual tokens among VQ-GAN-based Esser et al. (2021); Yu et al. (2021) and SEED-based Ge et al. (2023a, b). To check the effect of visual tokens only, we employ the same architecture and training configuration of Kim et al. (2024a) and only change the visual tokens. The performances of image captioning and image-to-speech captioning on COCO are shown in Table 10. In both tasks, SEED-based token achieves better performances. As SEED tokenizer is trained along with image-text association learning, its token achieves better performance in image-language association with fewer tokens (i.e., 784 vs. 32). Therefore, we employ SEED-based tokenizer for our experiments.
@@
Table 9: Ablation study to confirm the effects of vocabulary size of speech tokens on Image-to-Speech Captioning (BLEU, METEOR, ROUGE, CIDEr, SPICE) and Automatic Speech Recognition (WER).
Table 10: Captioning tasks’ performance comparisons using different visual token types.

### v2

@@ IV-B Metrics
Captioning tasks. For the image captioning and image-to-speech captioning tasks, we employ BLEU [87], METEOR [88], ROUGE [89], CIDEr [90], and SPICE [91]. BLEU, METEOR, ROUGE, and CIDEr measure n-gram overlaps between predictions and references. SPICE compares semantic propositional content. To measure these metrics for the speech outputs (i.e., image-to-speech captioning), we employ a pre-trained ASR model [60] to transcribe the speech into text following [13]. Then, the metric is measured on text-level. Across all metrics, higher value indicates better performance.
Image synthesis tasks. For the text- and speech-driven image synthesis tasks, we adopt the CLIP score [28] (i.e., the cosine similarity) to assess how similar the generated image is associated with the input text or speech, following [23, 92].
ASR and TTS. We employ the widely used Word Error Rate (WER) for ASR and TTS. For TTS, we transcribe the generated speech by using a pre-trained ASR model [60], following [63, 18]. Additionally, we employ Mean Opinion Score (MOS) and neural MOS to assess the quality of generated results.
TABLE IV: Image captioning performance (Image-to-Text Translation) comparisons on COCO and Flickr8k. We also report the performance of raw image-based models for reference purposes. “Num. MMT.” indicates the number of MMT tasks that the model is capable of.
@@
Raw Image SAT [26] 1 24.3 23.9 - - - 21.3 20.3 - - -
LaDic [93] 1 38.2 29.5 58.7 123.2 22.4 - - - - -
Kim et al[13, 27] 1 38.7 29.5 59.1 131.2 23.3 30.8 26.9 55.8 93.8 20.0
BLIP-2 [33] 1 38.5 32.5 60.1 135.8 26.4 30.4 29.4 56.4 99.6 22.5
Discrete Image Token Kim et al[13] 1 29.9 25.2 52.8 97.4 18.6 23.4 22.0 48.9 63.3 15.4
Ge et al[23, 24] 2 34.6 28.4 56.4 131.0 22.9 24.4 25.3 51.7 75.4 18.1
@@
TABLE V: Image-to-speech captioning performance (Image-to-Speech Translation) comparisons on COCO and Flickr8k. We also report the performance of raw image-based models for reference purposes. “Num. MMT.” indicates the number of MMT tasks that the model is capable of.
@@
Raw Image Wang et al[94] 1 - - - - - 3.5 11.3 23.2 8.0 -
Hsu et al[34] 1 23.3 21.2 47.8 73.2 14.9 12.5 14.5 39.1 24.5 9.5
Effendi et al[35] 1 - - - - - 14.8 17.4 32.9 45.8 -
Kim et al[13] 1 25.9 23.8 50.4 81.1 17.5 20.6 22.0 48.4 53.6 15.8
Discrete Image Token Kim et al[13] 1 20.1 21.4 46.4 64.0 15.0 16.7 19.6 44.2 41.2 13.1
@@
IV-C Implementation Details
The input image is resized into 224 $\times$ 224, while the output image size is 768 $\times$ 768. The input audio is resampled to 16kHz and tokenized using a pre-trained HuBERT base model [61]. The backbone model that models the multi-modal information is an encoder-decoder Transformer [71]. Both the encoder and decoder are initialized from a pre-trained BERT-Base configuration [50], unless otherwise specified. Therefore, each encoder and decoder contains 12 layers, where each layer has an embedding dimension of $768$, a feed-forward dimension of $3072$, and $12$ attention heads. If BERT-Large configuration is applied, it comprises 24 layers with an embedding dimension of $1024$, a feed-forward dimension of $4096$, and $16$ heads.
Each token is embedded through an embedding layer, where the total vocabulary size is $38914$ consisting of $8192$, $200$, and $30522$ vocabularies from image, speech, and text respectively. We add a modality type embedding that signifies which modality the input belongs to in the same manner as adding positional embeddings. We use speech samples up to a length of 384 tokens after deduplicating speech tokens. During the pre-training phase before applying BT, we use a batch size of 80 for each MMT task, a peak learning rate of $1e^{-4}$ with warm-up, Adam optimizer [95], and 350k steps. Then, with the intermediate pre-trained TMT model, BT is applied on ImageNet and CommonVoice as described in §$\S$§III-D. Then, we further train the model using the additional data for 250k steps. Finally, we finetune the model for 50k steps on COCO and Flickr8k with a peak learning rate of $1e^{-5}$. For the detokenization, we employ the decoder of SEED-2 [24] for image and a token-based HiFi-GAN vocoder [96, 97] trained on LJSpeech [98], for speech. To handle the removed duration information in the speech tokens during the deduplication, the vocoder is augmented with the duration predictor as in [99, 97]. Therefore, the speech token is repeated according to the predicted duration before generating the raw waveform. For decoding, we use beam search with a beam width of 5. Models are trained on 4 A40 (48GB) GPUs.
IV-D Results
Firstly, we show qualitative results for each task of TMT in Fig. 2. In the next sections, we compare single MMT models with the proposed TMT model. Then, we evaluate TMT against existing literature. We also conduct ablation studies on tokenizers, translation module, and the impact of using different modalities. Lastly, we include a human subjective evaluation on Image-to-Speech synthesis and TTS tasks.
IV-D1 Effectiveness of Unifying Tasks
@@
Table II compares the captioning performances, including image-to-text and image-to-speech translations. The results show that by unifying tasks, we can consistently obtain performance gain. In our analysis, this stems from the unified multi-modal encoder-decoder. The encoder-decoder can leverage complementary multi-modal information in learning language modeling, as both speech and text modalities contain linguistic information. Table III shows the comparison results for the remaining four tasks. Similar to captioning, the unified TMT model outperforms its single-task counterparts consistently, except for the text-to-image translation. However, the performance differences between single MMT and TMT in text-to-image translation are marginal, while the performance gain obtained by unifying tasks is substantial for speech-to-image translation. We conclude that unifying tasks is beneficial in terms of both practicality (i.e., by building a single model) and performance. More detailed analysis on different task combinations can be found in §$\S$§IV-D7.
IV-D2 Comparison with the literature
We compare TMT’s performance on each MMT task with the previous works.
@@ IV-D2 Comparison with the literature
Image Captioning Tasks. Table IV shows the results of image-to-text translation of TMT and other competing systems. We also report the performances of raw image-based systems for reference purposes. We observe that our TMT accomplishes comparable performances among state-of-the-art discrete image-based models. Notably, the proposed TMT is competitive with the LLM-based method [23] with a much smaller number of parameters. Note that while the modality translation module (excluding the tokenizer and detokenizer) of [23] is built on OPT 2.7B model [100], the proposed TMT employs a 270M multi-modal model. This shows the effectiveness of the proposed TMT compared to the method of finetuning LLM. Moreover, TMT can perform six MMT tasks while the previous methods can only perform two MMT tasks at maximum.
@@
Table V shows the direct image-to-speech captioning performances. TMT outperforms the recently proposed state-of-the-art [13] on both COCO and Flickr8k databases and even achieves comparable performances with raw-image based method. Please note that the vocoder used in [13] and TMT is the same. Since the decoder of the proposed TMT is trained using multi-tasks with text simultaneously, it better models linguistic information compared to the transfer learning from the text of [13].
@@
Image Synthesis Tasks. The upper part of Table VI displays the performances of text- and speech-driven image synthesis tasks. Following the literature, we evaluate the model on Flickr30k instead of using Flickr8k for the text-to-image synthesis tasks. The proposed TMT achieves the best CLIP similarity on both databases in text-to-image translation, indicating that TMT generates images that best describe the given text sentences. Please note that TMT still outperforms the competitors with more diverse MMT task abilities. Also, the speech-to-image translation performance of TMT on COCO is 68.88, close to that of text-to-image performance, while the cascaded system of [39] and Wav2Vec2.0 (W2V2) [60] achieves 62.87. We confirm that even if using speech directly as input, TMT correctly generates images described in the speech. We underline that this work is the first to explore high-resolution image synthesis at a resolution of 768$\times$768 from speech.
@@
ASR and TTS Tasks. The lower part of Table VI shows the performances of ASR and TTS. For evaluation purposes, we compare the ASR performances with pre-trained WavLM [62], HuBERT [61], and, W2V2 [60] models. The results show that TMT can correctly recognize the input speech into text by achieving lower WER compared to other pre-trained methods. For TTS, as the proposed model is a discrete token-based speech synthesis model, we compare the performances with UTUT [18] and the resynthesized speech (ReSyn) from the ground-truth speech tokens following [63, 18]. Please note that the vocoder used for ReSyn and TMT is the exact same model; the only difference between the two is how the input speech tokens for the vocoder are obtained. ReSyn employs the ground truth speech tokens, while TMT employs speech tokens that predicted by itself from the input text. Since the vocoder is exactly same, we can access how the speech tokens produced by TMT are in a good quality by comparing its performance with that of the resynthesis. Results demonstrate that even though the proposed TMT is trained using ground truth speech tokens (which inherently contain errors), it effectively learns the relationships between text and speech by suppressing undesired noises present in the ground truth speech tokens, and generates correct speech from input text.
TABLE VI: Text- and Speech-driven image synthesis, ASR, and TTS performance comparisons on COCO and Flickr8k. For TTS, we report the WER by feeding the synthesized audio into a pre-trained ASR model.
@@
IV-D3 Ablation on Speech and Image Tokens
@@
To confirm the effectiveness of the different multi-modal tokens, we perform ablation studies by differing configurations for speech and image tokens. First, we differentiate the number of k-means clustering of the speech tokenizer to obtain different token sizes. Then, we train single MMT models of image-to-speech captioning and ASR using 200, 500, and 1000 token sizes. Table VII shows the image-to-speech captioning and ASR performances. The results show that there is a trade-off in performances of ASR (WER) and image-to-speech captioning (BLEU, METEOR, ROUGE, CIDEr, SPICE). We confirm a consistent tendency where image-to-speech captioning performs better with a smaller number of speech tokens. In contrast, for ASR, employing a large number of tokens is beneficial. The trade-off in number of tokens can be understood as follows: since speech tokens are produced through quantization, the number of tokens is directly related to the level of information compression. Specifically, employing a higher number of speech tokens allows for the utilization of finer-grained speech features [63, 12]. However, the benefits of these finer-grained features depend on the task at hand. One possible intuition is that with a sufficient amount of speech tokens, the model can focus on generating speech with correct linguistic content, while an excessive number of similar-sounding tokens may cause distraction in content modeling. On the other hand, in ASR, utilizing a large number of speech tokens enables to capture subtle pronunciation differences. Since the performance gaps in ASR are smaller than those of image-to-speech captioning, we employ 200 speech tokens for all experiments.
Moreover, we evaluate the impact of using different SSL speech models as tokenizers on image-to-speech captioning and ASR tasks. To this end, we train the models with different tokens: one is tokenized using WavLM [62] following [12], and the other is tokenized using HuBERT following [63]. Similar to the previous experiment, we train single MMT models of image-to-speech captioning and ASR using each tokenizer. Table VIII shows the comparison between the two tokenizers. We note that using HuBERT as the speech tokenizer achieves better performances than using WavLM.
TABLE VII: Ablation study to confirm the effects of vocabulary size of speech tokens on COCO.
@@
Second, we investigate the different types of visual tokens among VQ-GAN-based [37, 103] and SEED-based [23, 24]. To check the effect of visual tokens only, we employ the same architecture and training configuration of [13] and only change the visual tokens. The performances of image captioning and image-to-speech captioning on COCO are shown in Table IX. In both tasks, SEED-based token achieves better performances. As SEED tokenizer is trained along with image-text association learning, its token achieves better performance in image-language association with fewer tokens (i.e., 784 vs. 32). Therefore, we employ SEED-based tokenizer for our experiments.
TABLE VIII: Performance comparisons using different SSL speech models as tokenizers on COCO.
  Token Type BLEU-4 METEOR ROUGE CIDEr SPICE WER
WavLM [62] 8.6 13.5 34.9 21.3 6.9 15.9
HuBERT [61] 22.5 21.5 47.1 69.1 15.0 6.9
TABLE IX: Captioning tasks’ performance comparisons using different visual token types on COCO.